18. What is the Discounted Payback Period for the project? Will you accept the project if the benchmark Discounted Payback Period is 3 years 8 months?
Answer #1

Payback period is the concept used in capital budgeting, this the time period required to recover the cost.

for example if the cost incurred is $100 and receive 50$ per year then the payback period is 2 years

in this concept we are not considering time value of money , that is the value of 1$ is more than the value of $1 tomorrow.

Discounted payback period : is the period of time required to recover cost incurred by consider the time value of money also

means all the future cash flows should be converted into present values by multiplying with present value interest factor ( PVIF table )

eg: initial cash outflow or cost incurred in the project is 100$ and the returns or inflows of cash is $50 per year and if discounted rate or cost of capital is 10%

for discounted payback period should calculate PV of the future cash flows

50/1.1 = 45.45 -----------> year 1

50/(1.1 * 1.1) = 41.32 -------> year 2

50/(1.1*1.1*1.1) =37.57 ----------> year 3

discounted payback period :  

$100- 45.45= 54.55 ------------1

$54.55- 41.32= $13.23----------2

$13.23 - 37.57= -24.34 so it is 2 years and some fraction of 3 year

=> 13.23/37.57 = 0.35 years

hence total discounting payback period is 2.35 years or 2 years 128 days

i think in the 2nd part of the question values are missing about the project

we can calculate the discounted payback period in the above manner of the particular project and if it takes less than the benchmark period that is 3 years and 8 months project can be accepted it means we are getting break even point earlier than the expected